So, This deals with perpendicular lines, Please show me examples of how to solve parallel, and perpendicular lines, and all the steps you used, not just the answer, I'll give a equation of each
Which equation has a graph parallel to the graph of 9x + 3y = -22
Which equation has a graph perpendicular to the graph of 7x + 14y - 8
Thx <3

Found 3 solutions by Boreal, Alan3354, solver91311:
Answer by Boreal(15235)

9x+3y=-22
slope intercept form
3y=-9x-22
divide by 3 to make y=1
y=-3x-22/3
any line of the form y=-3x + something will be parallel to this line unless it is y=-3x-22/3, in which case it is the same line.
parallel lines have the same slope.
====================================
the slopes of perpendicular lines have a product=-1
for the negative reciprocal, flip the fraction and change the sign.
the second one I will assume is written as 7x+14y=-8
14y=-7x-8
y=(-7x/14)-(8/14) =(-1/2)x-(8/14)
perpendicular line will have slope +2/1 or +2
It will have the form y=2x+ some number.

Answer by Alan3354(69443)

So, This deals with perpendicular lines, Please show me examples of how to solve parallel, and perpendicular lines, and all the steps you used, not just the answer, I'll give a equation of each
Which equation has a graph parallel to the graph of 9x + 3y = -22
Which equation has a graph perpendicular to the graph of 7x + 14y - 8
==================
Step 1, you find the slope of the given line.
To do that, solve the equation for y.
9x + 3y = -22
y = -3x - 22/3
That's the slope-intercept form.
The coefficient of x is the slope, m.
m = -3
====================
The slope of lines parallel have the same slope, -3
The slope of line perpendicular is the negative inverse.
1/3 for this example.
------------------------
that's all there is to it.

Answer by solver91311(24713)



Calculate the slope and the y-intercept of the line represented by your given equation. Then calculate the slopes and y-intercepts of the lines represented y the possible answers you are given. If the slopes are equal and the y-intercepts are different, then the lines are parallel. If the slopes are equal and the y-intercepts are the same, then the equations represent the same line. If the slopes are negative reciprocals of each other, for example, if one line has a slope of 3 and the other has a slope of , then the lines are perpendicular.

John

My calculator said it, I believe it, that settles it